OpenStudy (precal):

I need help with a limit limit as x approaches 0

OpenStudy (precal):

\[\frac{ \frac{ 1 }{ x+3 } -\frac{ 1 }{ 3 }}{ x }\]

OpenStudy (precal):

I haven't had a chance to do this problem at all. I forgot how to input this in wolframalpha.

ganeshie8 (ganeshie8):

thats same as the derivative of 1/(x+3) at x=0

OpenStudy (precal):

-1/9 ?

OpenStudy (precal):

You are so awesome :) You made my day :)

ganeshie8 (ganeshie8):

no, you're awesome :) \[f'(a) = \lim\limits_{x\to a}~\dfrac{f(x) - f(0)}{x-a}\] we plugin \(a=0, ~f(x) = \frac{1}{x+3}\)